OpenStudy (anonymous):

for each of the values of h given, when x is increased from 1 to 1+h, work out:

OpenStudy (anonymous):

(the change in x^2) / h h=1___ h=.1____ h=0.1_____ h=.001___

OpenStudy (anonymous):

can you write the change in x^2 in terms of h ?

OpenStudy (anonymous):

would it be (1+h-1)^2..?

OpenStudy (anonymous):

change in x^2 = (1+h)^2 - 1^2 = (1+h)^2 - 1

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so then now I can just plug in h values

OpenStudy (anonymous):

no first simplify

OpenStudy (anonymous):

it will be easier

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

this exercise should show you that as h approaches zero you get the value of the derivative of x^2 at x=1.

OpenStudy (anonymous):

managed to do it ?